# Estimating Knowledge in Large Language Models Without Generating a Single Token

### v1

@@ Abstract
To evaluate knowledge in large language models (LLMs), current methods query the model and then evaluate its generated responses. In this work, we ask whether evaluation can be done before the model has generated any text. Concretely, is it possible to estimate how knowledgeable a model is about a certain entity, only from its internal computation? We study this question with two tasks: given a subject entity, the goal is to predict (a) the ability of the model to answer common questions about the entity, and (b) the factuality of responses generated by the model about the entity. Experiments with a variety of LLMs show that KEEN, a simple probe trained over internal subject representations, succeeds at both tasks — strongly correlating with both the QA accuracy of the model per-subject and FActScore, a recent factuality metric in open-ended generation. Moreover, KEEN naturally aligns with the model’s hedging behavior and faithfully reflects changes in the model’s knowledge after fine-tuning. Lastly, we show a more interpretable yet equally performant variant of KEEN, which highlights a small set of tokens that correlates with the model’s lack of knowledge. Being simple and lightweight, KEEN can be leveraged to identify gaps and clusters of entity knowledge in LLMs, and guide decisions such as augmenting queries with retrieval.
@@ 1 Introduction
In this work, we take a step back and ask whether it is possible to evaluate the model’s knowledge before it generates any text, using only its internal computation. This view is analogous to human studies that show the effectiveness of non-verbal communication for assessing witness credibility in the courtroom Remland (1994); Denault et al. (2024). Concretely, we propose to evaluate how knowledgeable an LLM is about a given subject entity (e.g. Napoleon or Empire State Building), by considering only how it processes the name of that entity, and before it generates a single token.
@@
To tackle entity knowledge estimation, we capitalize on findings from recent interpretability works which show that, during inference, the hidden representations of an input entity capture many attributes related to it Geva et al. (2023); Meng et al. (2024), and often these attributes can be extracted with linear functions Hernandez et al. (2024). Therefore, we propose (§3) to estimate how knowledgeable a model is about a given entity by training simple probes, called KEEN (Knowledge Estimation of ENtities), over the model’s representations of the entity (Figure 1, upper left).
@@
We evaluate KEEN in two experimental settings (§4) of factual question answering (QA) and open-ended generation (OEG) of biographies. In the QA setting, we derive a set of questions per-subject for subjects in PopQA (Mallen et al., 2023) and evaluate how well KEEN predicts the model’s average accuracy per-subject across these questions. In the OEG setting, we evaluate the correlation of KEEN with FActScore (Min et al., 2023), a post-generation hallucination detector. In both settings and across models of different sizes and families — GPT2 Radford et al. (2019), Pythia Biderman et al. (2023), LLaMA2 Touvron et al. (2023), and Vicuna Chiang et al. (2023)KEEN consistently shows a strong correlation between 0.58-0.68 with model accuracy and 0.66-0.77 with factuality. Moreover, KEEN probes trained on entity representations show substantially stronger correlation with model accuracy and factuality than probes trained on commonly-used intrinsic features, such as fully-connected scores and self-attention activations, and external features, such as entity-popularity.
@@
Further analyzing the utility and features of KEEN5), we show that KEEN faithfully correlates with the model’s hedging behavior, i.e., the score predicted by KEEN decreases as the fraction of per-entity questions that a model hedges on increases. In addition, KEEN faithfully reflects changes in the model’s knowledge following fine-tuning: training LLaMA2 on Wikipedia articles about certain entities increases their KEEN score while scores for other entities tend to decrease. Lastly, we show that training KEEN on the vocabulary projections of entity representations nostalgebraist (2020); Geva et al. (2021) increases the probe’s interpretability without performance cost, identifying a small set of tokens that signal a lack of entity knowledge.
@@
To conclude, we present KEEN, a simple and lightweight approach for quantifying how knowledgeable a model is about a given entity from intrinsic properties, which well-estimates the accuracy and factuality of model outputs about the entity. We also show that KEEN scores are reflective of both hedging behavior and changes in entity-based knowledge over fine-tuning. KEEN could be used to inform developer decisions such as whether to augment queries with retrieval, discard certain queries (e.g. by abstaining), enhance models with external tools, or identify “holes” in the model’s knowledge to apply further training on. We release our code and data at https://github.com/dhgottesman/keen_estimating_knowledge_in_llms.
@@ 2 Entity Knowledge Estimation
Our goal is to evaluate how much knowledge an LLM captures about an entity from how it processes the entity’s name alone, without obtaining model responses and evaluating them post-generation. This view is motivated by growing evidence from interpretability works which find that, during model inference, knowledge is centralized in the hidden representations corresponding to named entities (Meng et al., 2024; Geva et al., 2023; Li et al., 2021).
@@
For a general information-seeking query $q$ about a subject $s$ (e.g. “Tell me facts about Napoleon” or “Generate a paragraph about Napoleon”), let =$\mathcal{R}=\{\langle c_{i},a_{i}\rangle\}_{i=1}^{m}$ be the set of claims in the response generated by $M$, each with a 0/1 label indicating its correctness with respect to external evidence. Claims can be extracted and evaluated for correctness using various automatic methods (e.g., Nenkova and Passonneau, 2004; Shapira et al., 2019; Zhang and Bansal, 2021). Given only the subject $s$, the task is to predict the portion of factually correct claims in $\mathcal{R}$, denoted as ${y^{(s)}_{OEG}:=\frac{1}{m}\sum_{i=1}^{m}a_{i}}$.
@@ 3 KEEN
Geva et al. (2023) showed that for a given subject in the input, LLMs construct an information-rich representation of the subject that encodes many of its attributes. Furthermore, subject attributes can be extracted from the subject representation with a simple linear function (Hernandez et al., 2024). We capitalize on these findings and propose to train a simple probe over the model’s representations of subjects to predict how much knowledge the model captures about them. In our following formulation (and the rest of the paper), we focus on widely-adopted transformer-based auto-regressive language models.
@@ 3.1 Features
We use the following sets of features for $\mathbf{z}^{(s)}$:
  • Hidden states (HS): We take the subject representation from multiple upper-intermediate layers, where attributes of the subject are often extracted during inference Geva et al. (2023); Meng et al. (2024) and are easier to disentangle Huang et al. (2024); Hernandez et al. (2024). To account for variations in the inference pass of different subjects, we choose 3 consecutive layers =$\mathcal{L}=\left\{\frac{3}{4}L+k\mid k\in\{-1,0,1\}\right\}$, from which we extract the hidden states ${\{\mathbf{h}^{(s)}_{\ell,s_{r}}\mid\ell\in\mathcal{L}\}}$. Then, we normalize these vectors (see details below) and average them into a $d$-dimensional feature vector.
@@ 4 Experiments
In this section, we evaluate KEEN and baselines that rely on different intrinsic and external features. We observed that the VP-50 probe obtained comparable performance while being significantly more interpretable (discussed in B.2) so we focus on evaluating the VP and VP-50 variants of this probe.
@@ 4.1 Experimental Setting
Examples for the two tasks are shown in Table 1. For both settings, we randomly split each dataset into disjoint sets of subjects: 65% train, 15% development, and 20% test. Importantly, the FActScore dataset and QA train set have a negligible number of overlapping subjects, 1 (0.2%), which allows us to test transfer learning between the two settings.
@@
  • Entity popularity (Pop.): It has been established that LLM performance is influenced by entity popularity Mallen et al. (2023); Kandpal et al. (2023); Yona et al. (2024). We follow previous works (e.g., Chen et al., 2021; Mallen et al., 2023; Cohen et al., 2024) and approximate entity popularity using statistics from Wikipedia. Concretely, we use the total number of monthly views of the entity’s page between the years 2000-2023.
@@
We analyze 7 auto-regressive language models across various sizes, latent spaces, and training objectives: GPT2-XL (Radford et al., 2019), Pythia 6B and 12B Biderman et al. (2023), LLaMA2 7B and 13B (Touvron et al., 2023), and Vicuna 13B (Chiang et al., 2023). [^3] We also analyzed Vicuna 7B, but due to its poor accuracy in the QA setting and inconsistent behavior, we omitted it in the main results. Results for Vicuna 7B can be found in §B. The vocabulary sizes range between 30K-50K tokens and the hidden state dimensions range from 4096-5120.
@@
For every model and subject $s$ in our data, we feed the model a generic prompt ‘‘This document describes [$s$]’’ and extract the features used for all methods: KEEN and the above baselines. Using these features, we obtain predictions for our two tasks for every method. For the Pop. baseline, we simply take the corresponding popularity value of the subject. We report Pearson correlation and the MSE between the predicted and gold scores, for every task, model and method. Correlation results are provided in §4.2 and the MSE results are reported in §B.
@@
Table 2: Correlation with the average QA accuracy for the KEEN QA probes and baselines.
@@ 4.2 Results
Table 2 and 3 show the QA and OEG results, respectively.
In both settings and across all models, KEEN probes trained on hidden representations and vocabulary projections demonstrate the strongest correlation of 0.60-0.68 with QA accuracy and 0.66-0.77 with FActScore. This shows that it is possible to predict how knowledgeable a model is about an entity from the entity’s hidden representations.
@@
Surprisingly, for the Pythia models even the KEEN OEG VP-50 probe strongly correlates with FActScore, indicating that there is a relatively small set of tokens which are influential in increasing/decreasing predicted accuracy. We further analyze these tokens in §5.4 and provide intuition for interpreting them. Moreover, we discuss the trade-off between interpretability and score correlation in B.2.
@@
Table 3: Correlation with FActScore for the KEEN OEG probes and baselines.
@@
Table 4: Transfer learning results, showing the correlation between FActScore and KEEN QA probes and baselines. Results are reported over all 500 subjects in the FActScore dataset.
@@
Since knowledge is centralized in the internal representations of entities, their use in estimating knowledge should transfer across different settings. Table 4 shows that the predictions of KEEN QA probes have a strong correlation of 0.60-0.62 with FActScore. Further, the correlation of KEEN QA probes with QA accuracy and FActScore are notably similar, e.g. 0.60 and 0.62 for Vicuna 13B KEEN QA HS probes, respectively. These results show that HS and VP features capture signals that generalize across settings, regardless of whether the task requires explicit (QA) or implicit (OEG) recall of factual knowledge by the model.
@@ 5.1 Correlation with Model Hedging
Figure 2 confirms this hypothesis, showing that the KEEN QA VP score decreases as the fraction of queries the model hedges on increases. This implies that models may hedge based on features of the model’s internal representations of the entity, similarly to KEEN.
@@ 5.2 Reflecting Changes in Model Knowledge
Figure 3: Changes in the KEEN QA score and average QA accuracy after fine-tuning LLaMA2 7B on paragraphs about a target subject, for target and non-target subjects. These results are aggregated over individually fine-tuning for 20 subjects.
Our experiments so far evaluated KEEN while keeping the underlying LLM fixed. A natural question that arises is whether changes in the model’s knowledge are reflected in changes in the KEEN score. We test this by fine-tuning LLaMA2-7B on paragraphs about a target subject and measuring changes in both the average QA accuracy and the KEEN QA score. Concretely, we sample 20 subjects from the QA test dataset and retrieve paragraphs from the Wikipedia page of each subject using BM25 Robertson et al. (1995). [^4] We use the Wikipedia dump from August 28, 2023. Then, we use LoRA Hu et al. (2022) to fine-tune $<0.5\%$ of the model’s parameters, separately for each subject. After fine-tuning for a certain target subject, we compute the KEEN score for that subject, as well as for 256 non-target subjects from the QA test dataset. The KEEN QA probe trained over the model’s hidden states before fine-tuning is used to compute these scores.
Figure 3 shows that on average, QA accuracy scores for the target entities increase by 0.16 and KEEN QA scores increase by 0.18, as models are fine-tuned on paragraphs related to them. Conversely, inline with works about catastrophic forgetting which find that models tend to forget information about entities observed in pre-training (Tirumala et al., 2022), the QA accuracy scores for non-target entities decrease after fine-tuning. However, the KEEN scores for non-target entities stay relatively constant. Since fine-tuning often doesn’t erase residual information in LLMs (Patil et al., 2024), and KEEN relies on intermediate representations, a possible explanation for this discrepancy is that information is still encoded in the representations but the model fails to recall it.

### v2

@@ Abstract
To evaluate knowledge in large language models (LLMs), current methods query the model and then evaluate its generated responses. In this work, we ask whether evaluation can be done before the model has generated any text. Concretely, is it possible to estimate how knowledgeable a model is about a certain entity, only from its internal computation? We study this question with two tasks: given a subject entity, the goal is to predict (a) the ability of the model to answer common questions about the entity, and (b) the factuality of open-ended responses generated by the model about the entity. Experiments with a variety of LLMs show that KEEN, a simple probe trained over internal subject representations, succeeds at both tasks — correlating with both the QA accuracy of the model per-subject and FActScore, a recent factuality metric in open-ended generation. Moreover, KEEN naturally aligns with the model’s hedging behavior and faithfully reflects changes in the model’s knowledge after fine-tuning. Lastly, we show a more interpretable yet equally performant variant of KEEN, which highlights a small set of tokens indicative of clusters and gaps in the model’s knowledge. Being simple and lightweight, KEEN can be leveraged to guide decisions such as when it is appropriate to apply further training or augment queries with retrieval.
@@ 1 Introduction
In this work, we take a step back and ask whether it is possible to evaluate the model’s knowledge before it generates any text, using only its internal computation. This view is analogous to human studies that show the effectiveness of assessing non-verbal communication for determining witness credibility in the courtroom Remland (1994); Denault et al. (2024). Concretely, we propose to evaluate how knowledgeable an LLM is about a given subject entity (e.g. Napoleon or Empire State Building), by considering only how it processes the name of that entity, and before it generates a single token.
@@
To tackle entity knowledge estimation, we capitalize on findings from recent interpretability works which show that, during inference, the hidden representations of an input entity capture many attributes related to it Geva et al. (2023); Meng et al. (2022), and often these attributes can be extracted with linear functions Hernandez et al. (2024b). Therefore, we propose (§3) to estimate how knowledgeable a model is about a given entity by training simple probes, called KEEN (Knowledge Estimation of ENtities), over the model’s representations of the entity (Figure 1, upper left).
@@
We evaluate KEEN in two experimental settings (§4) of factual question answering (QA) and open-ended generation (OEG) of biographies. In the QA setting, we derive a set of questions per-subject for subjects in PopQA (Mallen et al., 2023) and evaluate how well KEEN predicts the model’s average accuracy per-subject across these questions. In the OEG setting, we evaluate the correlation of KEEN with FActScore (Min et al., 2023), a post-generation hallucination detector. In both settings and across models of different sizes and families — GPT2 Radford et al. (2019), Pythia Biderman et al. (2023), LLaMA2 Touvron et al. (2023), and Vicuna Chiang et al. (2023)KEEN consistently shows correlation values between 0.58-0.68 with QA accuracy and 0.66-0.77 with OEG factuality. Moreover, KEEN probes trained on entity representations show substantially stronger correlation with both metrics than probes trained on commonly-used intrinsic features, such as fully-connected scores and self-attention activations, and external features, such as entity-popularity.
@@
Further analyzing the utility and features of KEEN5), we show that KEEN faithfully correlates with the model’s hedging behavior, i.e., the score predicted by KEEN decreases as the fraction of per-entity questions that a model hedges on increases. In addition, KEEN faithfully reflects changes in the model’s knowledge following fine-tuning: training LLaMA2 on Wikipedia articles about certain entities increases their KEEN score while scores for other entities tend to decrease. Lastly, we show that training KEEN on the vocabulary projections of entity representations nostalgebraist (2020); Geva et al. (2021) increases the probe’s interpretability without performance cost, identifying a small set of tokens that represent clusters or gaps in entity knowledge.
@@
To conclude, we present KEEN, a simple and lightweight approach for quantifying how knowledgeable a model is about a given entity from intrinsic properties, which well-estimates the accuracy and factuality of model responses about the entity. We also show that KEEN scores are reflective of both hedging behavior and changes in entity-based knowledge after fine-tuning. Practically, KEEN could be used to inform developer decisions such as whether to augment queries with retrieval, discard certain queries (e.g. by abstaining), enhance models with external tools, or identify “holes” in the model’s knowledge to apply further training on. We release our code and data at https://github.com/dhgottesman/keen_estimating_knowledge_in_llms.
@@ 2 Entity Knowledge Estimation
Our goal is to evaluate how much knowledge an LLM captures about an entity from how it processes the entity’s name alone, without obtaining model responses and evaluating them post-generation. This view is motivated by growing evidence from interpretability works which find that, during model inference, knowledge is centralized in the hidden representations corresponding to named entities (Meng et al., 2022; Geva et al., 2023; Li et al., 2021).
@@
For a general information-seeking query $q$ about a subject $s$ (e.g. “Tell me facts about Napoleon” or “Generate a paragraph about Napoleon”), let =$\mathcal{R}=\{\langle c_{i},a_{i}\rangle\}_{i=1}^{m}$ be the set of claims in the response generated by $M$, each with a 0/1 label indicating its correctness with respect to external evidence ($c_{i}$ denotes a claim and $a_{i}$ is its factuality label). Claims can be extracted and evaluated for correctness using various automatic methods (e.g., Nenkova and Passonneau, 2004; Shapira et al., 2019; Zhang and Bansal, 2021). Given only the subject $s$, the task is to predict the portion of factually correct claims in $\mathcal{R}$, denoted as ${y^{(s)}_{OEG}:=\frac{1}{m}\sum_{i=1}^{m}a_{i}}$.
@@ 3 KEEN
Geva et al. (2023) showed that for a given subject in the input, LLMs construct an information-rich representation of the subject that encodes many of its attributes. Furthermore, subject attributes can be extracted from the subject representation with a simple linear function (Hernandez et al., 2024b). We capitalize on these findings and propose to train a simple probe over the model’s representations of subjects to predict how much knowledge the model captures about them. In our following formulation (and the rest of the paper), we focus on widely-adopted transformer-based auto-regressive language models.
@@ 3.1 Features
We train different variants of KEEN probes, each taking as input one of the following sets of features for $\mathbf{z}^{(s)}$:
  • Hidden states (HS): We take the subject representation from multiple upper-intermediate layers, where attributes of the subject are often extracted during inference Geva et al. (2023); Meng et al. (2022) and are easier to disentangle Huang et al. (2024); Hernandez et al. (2024b). To account for variations in the inference pass of different subjects, we choose 3 consecutive layers =$\mathcal{L}=\left\{\frac{3}{4}L+k\mid k\in\{-1,0,1\}\right\}$, from which we extract the hidden states ${\{\mathbf{h}^{(s)}_{\ell,s_{r}}\mid\ell\in\mathcal{L}\}}$. Then, we normalize these vectors (see details below) and average them into a $d$-dimensional feature vector. A systematic evaluation supporting our choice of layers is presented in §A.2.
@@ 4 Experiments
In this section, we evaluate KEEN and baselines that rely on different intrinsic and external features. We observed that the VP-50 probe obtained comparable performance while being significantly more interpretable (discussed in §C.2) so we focus on evaluating the VP and VP-50 variants of this probe.
@@ 4.1 Experimental Setting
Examples for the two tasks are shown in Table 1. For both settings, we randomly split each dataset into disjoint sets of subjects: 65% train, 15% development, and 20% test. Importantly, the FActScore dataset and QA train set have a negligible number of overlapping subjects, 1 (0.2%), which allows us to test transfer learning between the two settings. In §B, we include additional details regarding dataset generation.
@@
  • Entity popularity (Pop.): It has been established that LLM performance is influenced by entity popularity Mallen et al. (2023); Kandpal et al. (2023); Yona et al. (2024). We follow previous works (e.g., Chen et al., 2021; Mallen et al., 2023; Cohen et al., 2024) and approximate entity popularity using statistics from Wikipedia. Concretely, we use the total number of monthly views of the entity’s page between the years 2000-2023.  [^3] We also computed thresholded log-popularity as implied by Figure 1 in Mallen et al. (2023). KEEN is superior across all settings and models, except Vicuna 13B in the OEG setting where correlation increased to 0.65 while KEEN achieved 0.66.
@@
We analyze 7 auto-regressive language models across various sizes, latent spaces, and training objectives: GPT2-XL (Radford et al., 2019), Pythia 6B and 12B Biderman et al. (2023), LLaMA2 7B and 13B (Touvron et al., 2023), and Vicuna 13B (Chiang et al., 2023). [^4] Vicuna 7B was also analyzed, but due to its poor accuracy in the QA setting and inconsistent behavior, we omitted it. The vocabulary sizes range between 30K-50K tokens and the hidden state dimensions range from 4096-5120.
@@
For every model and subject $s$ in our data, we feed the model a generic prompt ‘‘This document describes [$s$]’’ and extract the features used for all methods: KEEN and the above baselines. Using these features, we obtain predictions for our two tasks for every method. For the Pop. baseline, we simply take the corresponding popularity value of the subject. We report Pearson correlation, associated $p$-values ($p$), and the MSE between the predicted and gold scores, for every task, model and method. Correlation results are provided in §4.2, and the $p$-values and MSE results are reported in §C.
@@
Table 2: Correlation with average QA accuracy for KEEN QA probes and baselines on the QA test set.
@@ 4.2 Results
Tables 2 and 3 show QA and OEG results, respectively.
In both settings and across all models, KEEN probes trained on hidden representations and vocabulary projections demonstrate the strongest correlation of 0.60-0.68 with QA accuracy ($p\leq 3.43e^{-70}$) and 0.66-0.77 with FActScore ($p\leq 4.02e^{-6}$). This shows that it is possible to predict how knowledgeable a model is about an entity from the entity’s hidden representations.
@@
Surprisingly, for the Pythia models even the KEEN OEG VP-50 probe strongly correlates (Akoglu, 2018; Schober et al., 2018) with FActScore, indicating that there is a relatively small set of tokens which are influential in increasing and decreasing predicted accuracy. We further analyze these tokens in §5.4 and provide intuition for interpreting them. Moreover, we discuss the trade-off between interpretability and score correlation in §C.2.
@@
Table 3: Correlation with FActScore for KEEN OEG probes and baselines on the OEG test set.
@@
Table 4: Transfer learning results, showing the correlation between FActScore and KEEN QA probes and baselines. Results are reported over the full OEG dataset i.e. all 500 subjects in the unlabeled FActScore dataset.
@@
Since knowledge is centralized in the internal representations of entities, their use in estimating knowledge should transfer across different settings. Table 4 shows that the predictions of KEEN QA probes have a moderate to strong correlation (Akoglu, 2018; Schober et al., 2018) of 0.60-0.62 with FActScore ($p\leq 2.12e^{-5}$). Further, the correlation of KEEN QA probes with QA accuracy and FActScore are notably similar, e.g. 0.60 and 0.62 for Vicuna 13B KEEN QA HS probes, respectively. These results show that HS and VP features capture signals that generalize across settings, regardless of whether the task requires explicit (QA) or implicit (OEG) recall of factual knowledge by the model.
@@ 5.1 Correlation with Model Hedging
Figure 2 confirms this hypothesis, showing that the KEEN QA VP score decreases as the fraction of queries the model hedges on increases. This implies that models may hedge based on features of the model’s internal representations of the entity, similarly to KEEN. Further details regarding the choice of hedging phrases are provided in §B.3.
@@ 5.2 Reflecting Changes in Model Knowledge
Figure 3: Changes in the KEEN QA score and average QA accuracy after fine-tuning LLaMA2 7B on paragraphs about a target subject. These results are aggregated over individual fine-tuning processes for 20 target subjects.
Our experiments so far evaluated KEEN while keeping the underlying LLM fixed. A natural question that arises is whether changes in the model’s knowledge are reflected in changes in the KEEN score. We test this by fine-tuning LLaMA2 7B on paragraphs about a target subject and measuring changes in both the average QA accuracy and the KEEN QA score. Concretely, we sample 20 subjects from the QA test dataset and retrieve paragraphs from the Wikipedia page of each subject using BM25 Robertson et al. (1995). [^5] We use the Wikipedia dump from August 28, 2023. Then, we use LoRA Hu et al. (2022) to fine-tune $<0.5\%$ of the model’s parameters, separately for each subject. After fine-tuning for a certain target subject, we compute the KEEN score for that subject, as well as for 256 non-target subjects from the QA test dataset. The KEEN QA probe trained over the model’s hidden states before fine-tuning is used to compute these scores.
Figure 3 shows that on average, QA accuracy scores for the target entities increase by 0.16 and KEEN QA scores increase by 0.18 as models are fine-tuned on paragraphs related to them. Conversely, inline with works about catastrophic forgetting which find that models tend to forget information about entities observed in pre-training (Tirumala et al., 2022), the QA accuracy scores for non-target entities decrease after fine-tuning. However, the KEEN scores for non-target entities stay relatively constant. Since fine-tuning often doesn’t erase residual information in LLMs (Patil et al., 2024; Hong et al., 2024), and KEEN relies on intermediate representations, a possible explanation for this discrepancy is that information is still encoded in the representations but the model fails to recall it. To test this hypothesis, we selected 85 non-target entities whose decrease in QA accuracy after fine-tuning, $0.50\pm 0.18$ to $0.24\pm 0.15$, was not reflected in changes in their KEEN scores post fine-tuning, $0.46\pm 0.12$ to $0.49\pm 0.12$. We then use activation patching to recover QA accuracy from their intermediate fine-tuned representations, demonstrating that they still encode entity-related information (Ghandeharioun et al., 2024; Mosbach et al., 2020).
Skip fine-tuned layers to recover accuracy (FT subj.)
This experiment shows that information is recoverable from the fine-tuned representations when later layers in the fine-tuned model are bypassed. The procedure involves feeding the fine-tuned model a question, extracting only the subject representations (for all subject tokens) from an intermediate later, patching them into the penultimate layer of the fine-tuned model, and then measuring the patched QA accuracy per subject. Patching increases accuracy by $0.24\pm 0.18$ above fine-tuned accuracy (FT) and restores accuracy within $-0.01\pm 0.17$ of pre-trained accuracy (PT), per subject on average.
Apply pre-trained layers to recover accuracy (PT layer)
This experiment shows that information can still be extracted from fine-tuned representations using the pre-trained model. The procedure involves feeding the fine-tuned model a question, extracting all representations from an intermediate layer, patching them into the penultimate layer of the pre-trained model, and then measuring the patched QA accuracy per subject. Similarly, patching increases accuracy by $0.27\pm 0.19$ above FT and restores accuracy within $-0.03\pm 0.19$ of PT.
These experiments demonstrate that KEEN scores are reflective of the knowledge encoded in the intermediate representations, and that the estimation gap between KEEN scores and QA accuracy post fine-tuning is because the later layers in the fine-tuned model suppress the generation of this information. Further details about the patching procedure and results are presented in §C.3.